OpenStudy (jenniferjuice):

medal rewarded Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form.

OpenStudy (jenniferjuice):

Derive the equation of the parabola with a focus at (−2, 4) and a directrix of y = 6. Put the equation in standard form. f(x) = one fourthx2 − x + 4 f(x) = −one fourthx2 − x + 4 f(x) = one fourthx2 − x + 5 f(x) = −one fourthx2 − x + 5

OpenStudy (anonymous):

oh I think the answer is .. hold on

OpenStudy (anonymous):

I didn't try to solve it whatsoever

OpenStudy (jenniferjuice):

wow thanks